# Inference for multiple treatment effects using confounder importance learning

### v4

@@ 4.3. Abortion and Crime Study
To avoid confounding when estimating the association between abortion and crime, it is important to account for state and time effects for the period 1985 to 1997, and various other covariates. Donohue III and Levitt (2001) consider the log of prisoners per capita and of police per capita, unemployment and poverty rates, income and beer consumption per capita, the aid to families with dependent children (AFDC) program generosity, the existence of a concealed guns law, and the one-year lagged versions of these variables. Belloni et al. (2014) extend the analysis by adding quadratic covariate effects, interactions and linear and quadratic interactions with time. To account for state effects, they define the response as the increase in crime rates between two consecutive years, the treatment as the increase in abortion rates, and they include as covariates the within-state crime averages and the initial crime rates at 1985. They also force the inclusion of the year indicators in the model to avoid their estimates being driven by time dynamics in crime and abortion. The additions of Belloni et al. (2014) are done to reduce the misspecification of the outcome model, which could hamper causal interpretations. Our analysis keeps these covariates and overall we have $n=576$ observations, 1 treatment and $J=295$ covariates. Further details are provided in Section H.5.
@@
Table 1 summarizes the results obtained with different approaches, the previous ones and the one we develop in this article. The least-squares analysis of Donohue III and Levitt (2001) based on a pre-defined set of covariates find a statistically significant association between abortion and the three crime types. For comparison we also run a least-squares regression using all covariates considered by Belloni et al. (2014), which returns no statistically significant results and very wide confidence intervals. This is as expected in situations where covariates are strongly correlated with the treatment. The DL analysis of Belloni et al. (2014) also returns no statistically significant associations, which we update here using the latest version (0.3.1) of their R package hdm. The main difference between DL and Donohue III and Levitt (2001) is that the former selected numerous covariates that are associated to the treatment (abortion) but not to the outcome, i.e. likely instruments. For violent crime, a LASSO analysis of the outcome equation selected no covariates, whereas 9 covariates are selected in the abortion equation. DL then proceeded by regressing violent crime on those 9 covariates. As shown in Table 1, adding said covariates that are highly correlated with abortion causes a variance inflation in the estimated effect for the latter. Also note that there is little evidence that the covariates are needed in the outcome equation, e.g. only one of their naive $p$-values (not accounting for post-selection inference) are below 0.05 (Table S1). A similar situation occurs for property crime and murder. For property crime 13 covariates are selected (1 $p$-value below 0.05), and for murder it is 8 covariates (no $p$-value below $0.05$). Applying BAC to these data gives qualitatively similar results to DL, in that no significant treatment effect is detected. Again a potential issue is that many covariates have a non-negligible contribution, which can cause variance inflation, e.g. 98 covariates have marginal posterior inclusion probability above 0.5 for violent crime, 97 for property crime, and 80 for murder. ACPME did find a significant association for all three crime types. This is interesting because, as discussed, ACPME presents similarities to BAC but attempts to ameliorate variance inflation due to selecting instruments.
@@
We re-analyze the data of Belloni et al. (2014) with standard BMA (Beta-Binomial model prior) and our CIL methodology, also forcing the inclusion of the year indicators in the model, following Belloni et al. (2014). To explore sensitivity of the results to the prior, we obtain results under a default normal prior on the parameters with diagonal covariance and a MOM prior with default dispersion. As shown in Table 1, the results of our CIL approach lie somewhere in between the significant results found by Donohue III and Levitt (2001) and by ACPME, and the non-significant results found by DL, BAC and BMA. In the violent crime analysis, CIL finds moderate evidence for a negative association between abortion and crime. Under CIL all covariates have a negligible posterior inclusion probability. This is contrast to BMA where several covariates (6 for BMA normal, 2 for BMA MOM) have posterior inclusion probabilities above 0.1, and to DL which selects 8 covariates (which, as discussed, are likely to be instruments). We observe a similar situation for property crime, where CIL produces higher posterior probabilities for the existence of an association than its BMA counterpart. However, these are only moderate and the estimated effect is small. For murder CIL finds no evidence for an association with abortion. See Section H.5 for a discussion on the covariates selected by BMA and CIL in each analysis.
@@ 5. Discussion
The two main ingredients in our proposal are learning from data whether and to what extent covariate inclusion/exclusion should be encouraged to improve multiple treatment inference, and a convenient computational strategy to render the approach practical. By learning the relative importance of potential confounders one hopes to obtain a better balance between over-selection of instruments and under-selection of confounders.
These issues are practically relevant, e.g. in the salary data we show that one may underestimate the association black race and salary. Further, the proposed Bayesian framework naturally allows for posterior predictive inference on functions that depend on multiple parameters, such as the variation in salary jointly associated with multiple treatments. Interestingly, our analyses reveal little progress in the association between salary and potentially discriminatory factors such as gender or race in 2019 relative to 2010, nation- and state-wise. These results are conditional on covariates that include education, employment and other characteristics that affect salary. That is, our results reveal lower salary discrepancies in 2019 between races/genders, provided that two individuals have the same characteristics (and that they were hired in the first place). This analysis offers a complementary view to analyses that are unadjusted by confounders, and which may reveal equally interesting information. For example, if females migrated towards lower-paying occupational sections in 2019 and received a lower salary as a consequence, this would not be detected by our analysis, but would be revealed by an unadjusted analysis.
@@ 6. Acknowledgments
David Rossell gratefully acknowledges support from grant Consolidación investigadora CNS2022-135963 by the AEI (Government of Spain), Ayudas Fundación BBVA Proyectos de Investigación Científica en Matemáticas 2021, Europa Excelencia EUR2020-112096 from the AEI/10.13039/501100011033 and European Union NextGenerationEU/PRT, and grant PID2022-138268NB-I00 financed by MCIN/AEI/10.13039/501100011033 and the FSE+.
Appendix A Proof of Proposition 3.1
The empirical Bayes estimate writes
For short, denote $H({\bm{\theta}})=p({\mathbf{y}}\mid{\bm{\theta}})$ and $h_{j}({\bm{\theta}})=p_{j}(\gamma_{j}\mid{\bm{\theta}})$, where generically $\nabla_{{\bm{\theta}}}\log H({\bm{\theta}})=\nabla_{{\bm{\theta}}}H({\bm{\theta}})/H({\bm{\theta}})$. Under the assumptions of Proposition 3.1
(16) $\nabla_{{\bm{\theta}}}H({\bm{\theta}})=\sum_{({\bm{\gamma}},{\bm{\delta}})}p({\mathbf{y}}\mid{\bm{\gamma}},{\bm{\delta}})p({\bm{\delta}})\nabla_{{\bm{\theta}}}\prod_{j=1}^{J}h_{j}({\bm{\theta}})$
$=$ $\sum_{({\bm{\gamma}},{\bm{\delta}})}p({\mathbf{y}}\mid{\bm{\gamma}},{\bm{\delta}})p({\bm{\delta}})\sum_{j=1}^{J}\left(\nabla_{{\bm{\theta}}}h_{j}({\bm{\theta}})\prod_{j\neq l}h_{l}({\bm{\theta}})\right).$
Denoting $\mathbf{f}_{j}=(1,f_{j,1},\dots,f_{j,T})^{\intercal}$, direct algebra gives
(17) $\nabla_{{\bm{\theta}}}h_{j}({\bm{\theta}})=(1-2\rho)\mathbf{f}_{j}(\gamma_{j}-\pi_{j}({\bm{\theta}}))h_{j}({\bm{\theta}}).$
Then, replacing (17) into (16)
$\nabla_{{\bm{\theta}}}H({\bm{\theta}})=\sum_{({\bm{\gamma}},{\bm{\delta}})}p({\mathbf{y}}\mid{\bm{\gamma}},{\bm{\delta}})p({\bm{\delta}})\sum_{j=1}^{J}(1-2\rho)\mathbf{f}_{j}(\gamma_{j}-\pi_{j}({\bm{\theta}}))\prod_{l=1}^{J}h_{l}({\bm{\theta}})\\ =\sum_{j=1}^{J}(1-2\rho)\mathbf{f}_{j}\sum_{({\bm{\gamma}},{\bm{\delta}})}(\gamma_{j}-\pi_{j}({\bm{\theta}}))p({\mathbf{y}}\mid{\bm{\gamma}},{\bm{\delta}})p({\bm{\delta}},{\bm{\gamma}}\mid{\bm{\theta}})\\ =\sum_{j=1}^{J}(1-2\rho)\mathbf{f}_{j}\left[(1-\pi_{j}({\bm{\theta}}))\sum_{({\bm{\gamma}},{\bm{\delta}}):\gamma_{j}=1}p({\mathbf{y}},{\bm{\delta}},{\bm{\gamma}}\mid{\bm{\theta}})\right.\left.-\pi_{j}({\bm{\theta}})\sum_{({\bm{\gamma}},{\bm{\delta}}):\gamma_{j}=0}p({\mathbf{y}},{\bm{\delta}},{\bm{\gamma}}\mid{\bm{\theta}})\right].$
Finally
$\nabla_{{\bm{\theta}}}\log H({\bm{\theta}})=\frac{\nabla_{{\bm{\theta}}}H({\bm{\theta}})}{H({\bm{\theta}})}=\\ \sum_{j=1}^{J}(1-2\rho)\mathbf{f}_{j}\left[(1-\pi_{j}({\bm{\theta}}))\frac{\sum_{({\bm{\gamma}},{\bm{\delta}}):\gamma_{j}=1}p({\mathbf{y}},{\bm{\delta}},{\bm{\gamma}}\mid{\bm{\theta}})}{\sum_{({\bm{\gamma}},{\bm{\delta}})}p({\mathbf{y}},{\bm{\delta}},{\bm{\gamma}}\mid{\bm{\theta}})}-\right.\left.\pi_{j}({\bm{\theta}})\frac{\sum_{({\bm{\gamma}},{\bm{\delta}}):\gamma_{j}=0}p({\mathbf{y}},{\bm{\delta}},{\bm{\gamma}}\mid{\bm{\theta}})}{\sum_{({\bm{\gamma}},{\bm{\delta}})}p({\mathbf{y}},{\bm{\delta}},{\bm{\gamma}}\mid{\bm{\theta}})}\right]\\ =\sum_{j=1}^{J}(1-2\rho)\mathbf{f}_{j}[(1-\pi_{j}({\bm{\theta}}))P(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}})-\pi_{j}({\bm{\theta}})(1-P(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}}))]\\ =(1-2\rho)\sum_{j=1}^{J}\mathbf{f}_{j}\left[P(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}})-\pi_{j}({\bm{\theta}})\right].$
$\blacksquare$
Appendix B Proof of Proposition 3.2
Consider the optima of the marginal likelihood,
Optimizing this expression yields
Appendix C Supplementary material
Appendix D Over-selection bias
D.1. Discussion
@@ D.2. Derivations
Appendix E Derivations
E.1. Derivation of Equation (14)

### v6

@@ 5.3. Abortion and Crime Study
To avoid confounding when estimating the association between abortion and crime, it is important to account for state and time effects for the period 1985 to 1997, and various other covariates. Donohue III and Levitt (2001) consider the log of prisoners per capita and of police per capita, unemployment and poverty rates, income and beer consumption per capita, the aid to families with dependent children (AFDC) program generosity, the existence of a concealed guns law, and the one-year lagged versions of these variables. Belloni et al. (2014) extend the analysis by adding quadratic covariate effects, interactions and linear and quadratic interactions with time. To account for state effects, they define the outcome as the increase in crime rates between two consecutive years, the treatment as the increase in abortion rates, and they include as covariates the within-state crime averages and the initial crime rates at 1985. They also force the inclusion of the year indicators in the model to avoid their estimates being driven by time dynamics in crime and abortion. The additions of Belloni et al. (2014) are done to reduce the misspecification of the outcome model, which could hamper causal interpretations. Our analysis keeps these covariates and overall we have $n=576$ observations, 1 treatment and $J=295$ covariates. See Section S8.5 for further details.
@@
Table 1 summarizes the results obtained with different approaches, the previous ones and the one we develop in this article. The least-squares analysis of Donohue III and Levitt (2001) based on a pre-defined set of covariates find a statistically significant association between abortion and the three crime types. For comparison we also run a least-squares regression using all covariates considered by Belloni et al. (2014), which returns no statistically significant results and very wide confidence intervals. This is as expected in situations where covariates are strongly correlated with the treatment. The DL analysis of Belloni et al. (2014) also returns no statistically significant associations, using the latest version (0.3.1) of their R package hdm. The main difference between DL and Donohue III and Levitt (2001) is that the former selected numerous covariates that are associated to the treatment (abortion) but not to the outcome, i.e. likely instruments. For violent crime, a LASSO analysis of the outcome equation selected no covariates, whereas 9 covariates are selected in the abortion equation. DL then proceeds by regressing violent crime on those 9 covariates. As shown in Table 1, adding said covariates that are highly correlated with abortion causes a variance inflation in the estimated effect for the latter. Also note that there is little evidence that the covariates are needed in the outcome equation, e.g. only one of their naive $p$-values (not accounting for post-selection inference) are below 0.05 (Table S1). A similar situation occurs for property crime and murder. For property crime 13 covariates are selected (1 $p$-value below 0.05), and for murder it is 8 covariates (no $p$-value below $0.05$). Applying BAC to these data gives qualitatively similar results to DL, in that no significant treatment effect is detected. Again a potential issue is that many covariates have a non-negligible contribution, which can cause variance inflation, e.g. 98 covariates have marginal posterior inclusion probability above 0.5 for violent crime, 97 for property crime, and 80 for murder. ACPME did find a significant association for all three crime types. This is interesting because, as discussed, ACPME presents similarities to BAC but attempts to ameliorate variance inflation due to selecting instruments.
@@
We re-analyze the data of Belloni et al. (2014) with standard BMA (Beta-Binomial model prior) and our CIL methodology, also forcing the inclusion of the year indicators in the model, following Belloni et al. (2014). To explore sensitivity of the results to the prior, we obtain results under a default normal prior on the parameters with diagonal covariance and a MOM prior with default dispersion. As shown in Table 1, the results of our CIL approach lie somewhere in between the significant results found by Donohue III and Levitt (2001) and by ACPME, and the non-significant results found by DL, BAC and BMA. In the violent crime analysis, CIL finds moderate evidence for a negative association between abortion and crime. Under CIL all covariates have a negligible posterior inclusion probability. This is contrast to BMA where several covariates (6 for BMA normal, 2 for BMA MOM) have posterior inclusion probabilities above 0.1, and to DL which selects 8 covariates (which, as discussed, are likely to be instruments). We observe a similar situation for property crime, where CIL produces higher posterior probabilities for the existence of an association than its BMA counterpart. However, these are only moderate and the estimated effect is small. For murder CIL finds no evidence for an association with abortion. See Section S8.5 for a discussion on the covariates selected by BMA and CIL in each analysis.
@@ 6. Discussion
The two main ingredients in our proposal are learning from data whether and to what extent covariate inclusion/exclusion should be encouraged to improve multiple treatment inference, and a convenient computational strategy to render the approach practical. Our framework learns from data whether the data-generating truth is of a high, neutral or low confounding nature, as measured by our novel confounding coefficient. One then hopes to obtain a better balance between over-selection of instruments and under-selection of confounders.
These issues are practically relevant, e.g. in the salary data we show that one may underestimate the association black race and salary. Further, the proposed Bayesian framework naturally allows for posterior predictive inference on functions that depend on multiple parameters, such as the variation in salary jointly associated with multiple treatments. Interestingly, our analyses reveal little progress in the association between salary and potentially discriminatory factors such as sex or race in 2019 relative to 2010, nation- and state-wise. These results are conditional on covariates that include education, employment and other characteristics that affect salary. That is, our results reveal lower salary discrepancies in 2019 between races/sex, provided that two individuals have the same characteristics (and that they were hired in the first place). This analysis offers a complementary view to analyses that are unadjusted by confounders, and which may reveal equally interesting information. For example, if females migrated towards lower-paying occupational sections in 2019 and received a lower salary as a consequence, this would not be detected by our analysis, but would be revealed by an unadjusted analysis.
@@ 7. Acknowledgments
DR gratefully acknowledges support from grant Consolidación investigadora CNS2022-135963 by the AEI (Government of Spain), Ayudas Fundación BBVA Proyectos de Investigación Científica en Matemáticas 2021, Europa Excelencia EUR2020-112096 from the AEI/10.13039/501100011033 and European Union NextGenerationEU/PRT, and grant PID2022-138268NB-I00 financed by MCIN/AEI/10.13039/501100011033 and the FSE+.
Supplementary material
8. Over-selection bias
8.1. Discussion
@@ 8.2. Derivations
9. Proof of Proposition 3.1
We first state and prove an auxiliary result regarding the gradient of $p(\gamma_{j}\mid{\bm{\theta}})$, and subsequently prove the proposition. To ease notation, let $\mathbf{f}_{j}=(1,f_{j,1},\dots,f_{j,T})^{\intercal}$ be the vector of features for covariate $j$, including the intercept.
9.1. Auxiliary result
Recall that the prior inclusion probability for covariate $j$ is
$\pi_{j}({\bm{\theta}})=\begin{cases}\bar{\rho}\mbox{, if }\tilde{\pi}_{j}({\bm{\theta}})\leq\bar{\rho}\\ \tilde{\pi}_{j}({\bm{\theta}})\mbox{, if }\tilde{\pi}_{j}({\bm{\theta}})\in(\underline{\rho},\bar{\rho})\\ \bar{\rho}\mbox{, if }\tilde{\pi}_{j}({\bm{\theta}})\geq\bar{\rho}\end{cases}$
where $\tilde{\pi}_{j}({\bm{\theta}})=(1+e^{-{\bf f}_{j}^{\top}{\bm{\theta}}})^{-1}$.
Let $p(\gamma_{j}\mid{\bm{\theta}})=\pi_{j}({\bm{\theta}})^{\gamma_{j}}[1-\pi_{j}({\bm{\theta}})]^{1-\gamma_{j}}$ be the corresponding prior probability mass function. We prove that
(19) $\nabla_{\bm{\theta}}p(\gamma_{j}\mid{\bm{\theta}})=\begin{cases}0\mbox{, if }\tilde{\pi}_{j}({\bm{\theta}})<\underline{\rho}\mbox{ or }\tilde{\pi}_{j}({\bm{\theta}})>\bar{\rho}\\ \mbox{undefined, if }\tilde{\pi}_{j}({\bm{\theta}})=\underline{\rho}\mbox{ or }\tilde{\pi}_{j}({\bm{\theta}})=\bar{\rho}\\ {\bf f}_{j}p(\gamma_{j}\mid{\bm{\theta}})[\gamma_{j}-\pi_{j}(\theta)]\mbox{, if }\tilde{\pi}_{j}({\bm{\theta}})\in(\underline{\rho},\bar{\rho})\end{cases}$
The first line in (19) holds trivially, since in that case $p(\gamma_{j}\mid{\bm{\theta}})$ does not depend on ${\bm{\theta}}$. The second line in (19) follows immediately by proving the third line, since then the directional derivatives at $\underline{\rho}$ and $\bar{\rho}$ do not match. To prove the third line in (19), note that when $\tilde{\pi}_{j}({\bm{\theta}})\in(\underline{\rho},\bar{\rho})$ we have
$\nabla_{\bm{\theta}}\pi_{j}({\bm{\theta}})=\nabla_{\bm{\theta}}\tilde{\pi}_{j}({\bm{\theta}})=\frac{{\bf f}_{j}e^{-{\bf f}_{j}^{\top}{\bm{\theta}}}}{(1+e^{-{\bf f}_{j}^{\top}{\bm{\theta}}})^{2}}={\bf f}_{j}\pi_{j}({\bm{\theta}})[1-\pi_{j}({\bm{\theta}})].$
Finally, we obtain $\nabla_{\bm{\theta}}p(\gamma_{j}\mid{\bm{\theta}})$ separately for the $\gamma_{j}=1$ and $\gamma_{j}=0$ cases. For the case $\gamma_{j}=1$,
$\nabla_{\bm{\theta}}p(\gamma_{j}=1\mid{\bm{\theta}})=\nabla_{\bm{\theta}}\pi_{j}({\bm{\theta}})={\bf f}_{j}\pi_{j}({\bm{\theta}})[\gamma_{j}-\pi_{j}({\bm{\theta}})]={\bf f}_{j}p(\gamma_{j}\mid{\bm{\theta}})[\gamma_{j}-\pi_{j}({\bm{\theta}})],$
proving the result. For the case $\gamma_{j}=0$,
$\nabla_{\bm{\theta}}p(\gamma_{j}=0\mid{\bm{\theta}})=\nabla_{\bm{\theta}}[1-\pi_{j}({\bm{\theta}})]=-{\bf f}_{j}\pi_{j}({\bm{\theta}})[1-\pi_{j}({\bm{\theta}})]={\bf f}_{j}p(\gamma_{j}\mid{\bm{\theta}})[\gamma_{j}-\pi_{j}({\bm{\theta}})],$
since $p(\gamma_{j}=0\mid{\bm{\theta}})=1-\pi_{j}({\bm{\theta}})$, again proving the result.
9.2. Proof of Proposition 3.1
The empirical Bayes estimate writes
For short, denote $H({\bm{\theta}})=p({\mathbf{y}}\mid{\bm{\theta}})$ where generically $\nabla_{{\bm{\theta}}}\log H({\bm{\theta}})=\nabla_{{\bm{\theta}}}H({\bm{\theta}})/H({\bm{\theta}})$. Under the assumptions of Proposition 3.1
(20) $\nabla_{{\bm{\theta}}}H({\bm{\theta}})=\sum_{({\bm{\gamma}},{\bm{\delta}})}p({\mathbf{y}}\mid{\bm{\gamma}},{\bm{\delta}})p({\bm{\delta}})\nabla_{{\bm{\theta}}}\prod_{j=1}^{J}p(\gamma_{j}\mid{\bm{\theta}})$
$=$ $\sum_{({\bm{\gamma}},{\bm{\delta}})}p({\mathbf{y}}\mid{\bm{\gamma}},{\bm{\delta}})p({\bm{\delta}})\sum_{j=1}^{J}\left(\nabla_{{\bm{\theta}}}p(\gamma_{j}\mid{\bm{\theta}})\prod_{j\neq l}p(\gamma_{l}\mid{\bm{\theta}})\right).$
Replacing (19) into (20)
$\nabla_{{\bm{\theta}}}H({\bm{\theta}})=\sum_{({\bm{\gamma}},{\bm{\delta}})}p({\mathbf{y}}\mid{\bm{\gamma}},{\bm{\delta}})p({\bm{\delta}})\sum_{j:\pi_{j}({\bm{\theta}})\in(\underline{\rho},\bar{\rho})}\mathbf{f}_{j}(\gamma_{j}-\pi_{j}({\bm{\theta}}))\prod_{l=1}^{J}h_{l}({\bm{\theta}})\\ =\sum_{j:\pi_{j}({\bm{\theta}})\in(\underline{\rho},\bar{\rho})}\mathbf{f}_{j}\sum_{({\bm{\gamma}},{\bm{\delta}})}(\gamma_{j}-\pi_{j}({\bm{\theta}}))p({\mathbf{y}}\mid{\bm{\gamma}},{\bm{\delta}})p({\bm{\delta}},{\bm{\gamma}}\mid{\bm{\theta}})\\ =\sum_{j:\pi_{j}({\bm{\theta}})\in(\underline{\rho},\bar{\rho})}\mathbf{f}_{j}\left[(1-\pi_{j}({\bm{\theta}}))\sum_{({\bm{\gamma}},{\bm{\delta}}):\gamma_{j}=1}p({\mathbf{y}},{\bm{\delta}},{\bm{\gamma}}\mid{\bm{\theta}})\right.\left.-\pi_{j}({\bm{\theta}})\sum_{({\bm{\gamma}},{\bm{\delta}}):\gamma_{j}=0}p({\mathbf{y}},{\bm{\delta}},{\bm{\gamma}}\mid{\bm{\theta}})\right].$
Finally
$\nabla_{{\bm{\theta}}}\log H({\bm{\theta}})=\frac{\nabla_{{\bm{\theta}}}H({\bm{\theta}})}{H({\bm{\theta}})}=\\ \sum_{j:\pi_{j}({\bm{\theta}})\in(\underline{\rho},\bar{\rho})}\mathbf{f}_{j}\left[(1-\pi_{j}({\bm{\theta}}))\frac{\sum_{({\bm{\gamma}},{\bm{\delta}}):\gamma_{j}=1}p({\mathbf{y}},{\bm{\delta}},{\bm{\gamma}}\mid{\bm{\theta}})}{\sum_{({\bm{\gamma}},{\bm{\delta}})}p({\mathbf{y}},{\bm{\delta}},{\bm{\gamma}}\mid{\bm{\theta}})}-\right.\left.\pi_{j}({\bm{\theta}})\frac{\sum_{({\bm{\gamma}},{\bm{\delta}}):\gamma_{j}=0}p({\mathbf{y}},{\bm{\delta}},{\bm{\gamma}}\mid{\bm{\theta}})}{\sum_{({\bm{\gamma}},{\bm{\delta}})}p({\mathbf{y}},{\bm{\delta}},{\bm{\gamma}}\mid{\bm{\theta}})}\right]\\ =\sum_{j:\pi_{j}({\bm{\theta}})\in(\underline{\rho},\bar{\rho})}\mathbf{f}_{j}[(1-\pi_{j}({\bm{\theta}}))P(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}})-\pi_{j}({\bm{\theta}})(1-P(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}}))]\\ =\sum_{j:\pi_{j}({\bm{\theta}})\in(\underline{\rho},\bar{\rho})}\mathbf{f}_{j}\left[P(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}})-\pi_{j}({\bm{\theta}})\right].$
$\blacksquare$
10. Proof of Proposition 3.2
Consider the optima of the marginal likelihood,
Optimizing this expression yields
11. Derivations
11.1. Derivation of Equation 3.9